OpenStudy (anonymous):

Prime Factors of 125-y3.

OpenStudy (lgbasallote):

we have a Edit Question button you know :)

OpenStudy (lgbasallote):

anyway..use the formula for difference of two cubes here \[\large a^3 - b^3 = (a-b)(a^2 + ab + b^2)\]

OpenStudy (anonymous):

\[(5^3 - y^3)\] Use the above formula to solve it...

OpenStudy (anonymous):

How about the prime factors of \[x ^{3}-1\] ?

OpenStudy (anonymous):

Just same.. You will use the same formula for this.. \[(x^3 - 1^3) = ???\] Use the above formula to find it..

OpenStudy (anonymous):

what is the prime actors of 8-\[x ^{3}\] ?

OpenStudy (lgbasallote):

still same.. 8 is 2^3

OpenStudy (lgbasallote):

x^3 is x^3

OpenStudy (anonymous):

Using same formula : \[2^3 - x^3\]

OpenStudy (anonymous):

aa .. ThankYou ! :]

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

how about the product o monomial by a binomial ? 3(x+5) ?

OpenStudy (anonymous):

Just distribute 3 inside the brackets: \[a(b+c) = ab + ac\]

OpenStudy (anonymous):

is my answer right ? 3x+15 ?

OpenStudy (anonymous):

Yes you are right...

OpenStudy (anonymous):

yey :]] and how about 4x(x-3) = 4x\[^{2}\]-12x ? is that's right ?

OpenStudy (anonymous):

Yes have faith on yourself you are very much correct in doing this..

OpenStudy (lgbasallote):

heh good one @eiznehr.12 :D

OpenStudy (anonymous):

hehe :]] Thankyou .. i just multiplied the monomial to the terms inside the parenthesis ? :]

OpenStudy (anonymous):

Yes that is what you are supposed to do...

OpenStudy (anonymous):

but I don't know how to factorize \[x ^{3}\]-8 ..

OpenStudy (lgbasallote):

difference of two cubes still...x^3 is x^3 and 8 is 2^3

OpenStudy (anonymous):

\[(x^3 - 8) = (x^3 - 2^3)\] Here, a = x and b = 2 Put these values of a and b in the formula written above and try it once on your own..

OpenStudy (anonymous):

hmmmm .. i got (x-2)(y2+2x+4)

OpenStudy (anonymous):

How come y^2??

OpenStudy (anonymous):

You almost did that right but one mistake only...

OpenStudy (anonymous):

sorry it's x :]

OpenStudy (anonymous):

hehe

OpenStudy (anonymous):

x or x^2???

OpenStudy (anonymous):

its x^2

OpenStudy (anonymous):

Yes now you are right: \[x^3 - 8 = x^3 - 2^3 = (x-2)(x^2 + 2x + 4)\]

OpenStudy (anonymous):

yes ! :]] ThanksALot :]]

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

ThankYou again :] hehe

OpenStudy (anonymous):

Welcome again...